Fermi National Laboratory


Accelerator Update

Wednesday October 1 through Wednesday October 3
All shifts began in shutdown.

On Wednesday technicians performed various computer and electrical work on the accelerators. CHL was without power during the day due to high voltage maintenance work.

On Thursday, techs worked on the Tevatron (TeV) air compressors. Water techs began filling the Linac header from CUB.

TEVATRON

  • The original installation of 59 magnet stands in E, B and A sectors has been completed. Another 15 dipole magnet stands have been added and will be changed next week.
  • All smart bolt work has been completed. Technical Division (TD) people are analyzing lift data and will re-measure when the TeV is cold.
  • Conning tower work at A-4 is done. The TD crew is moving onto A-2.
  • The new bypass at A-48 was installed. However, the single-phase connections had to be cut off and replaced on Thursday. The bypass will be re-installed today. Interfacing will happen next week.
  • All the Lambertsons have had liners installed and a survey completed. Vacuum assembly continues. Techs will survey the BPM and strip lines this morning. The Lambertsons should be under vacuum by Saturday.
  • Techs have begun installing the HLS system piping. Techs will setup a two-sensor test section today and then test the sensor heads. Techs continue to plumb the water pipe around the spool pieces. We found a broken ion pump at B-3 that hadn't been pumping for 2 weeks. A vacuum pump out is now underway.
  • TeV Net laser tracker crews have finished with A, B, C, D, and F sectors; a small part of E sector remains to be done.
  • The E-2 cold leak work continues. Techs replaced a bellows on a spool piece at E-26. The entire system is under vacuum.
  • Experts have completed checking the roll measurements data; they are rechecking a few sections. Crews finished calibrating the protractors and tilt meters last weekend.
  • Crews continue to pull cable for new octupole circuits. BPM testing is approximately 60% done.
  • A crew has finished the core drilling for 6 new sump pits.
  • The Collider Detector Facility (CDF) wall has been built in the Tevatron. The steel work inside of CDF will be completed in 2 weeks. We will monitor the low beta quads with tilt meters during the steel stacking. There will be a final survey before startup
  • Over all, our vacuum work is 1 week behind schedule.

SWITCHYARD
The Switchyard experts hope to finish the following work on the SY120 beam line and associated elements during this shutdown.

  • The Main Injector (MI) 52 septa are being realigned since it appears that replacing the motors caused an alignment problem.
  • Certain collimators in Enclosure-B are being removed in order to remove an aperture restriction.
  • The two remaining septa in Enclosure-B are being removed from the beam line since we plan on using only the triple split for beam splitting purposes.
  • Two new multiwire chambers are replacing the SWICs at S94 and S208. These can be left in the beam during normal SY120 running and will thus be a 'comfort' display for the MCR and the experimenters.

Startup will follow the same pattern as before the shutdown with engineering tests in the MTest beam line and the start of engineering tests in the MCenter beam line after the MCenter assessment is approved.

LINAC
The Linac low-level water system upgrade is 75% complete:

  • Cavity work, 72% complete
  • RF work, 73% complete
  • Valve work, 90% complete

    RECYCLER

    • 11 of 27 Recycler ring bake outs have been completed: 41% complete
    • 16 of 18 gradient magnet heater tape repairs have been finished: 89% complete
    • 82 of 100 quadrupole heater tape repairs have been finished: 82% complete
    • 5 of 6 ion pumps have been upgraded: 83% complete
    • 9 of 11 additional vacuum readback installations have been finished: 82% complete
    • 2 of 4 of the LEP corrector magnets have been replaced: 50% complete
    • Rework of the 620 region has yet to begin

    Accelerator Update Archive

    More Information
    For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

    Comments and Suggestions
    What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



  • last modified 10/3/2003   email Fermilab